ISOLDE Technical Coordination Meeting (LS3)
-
-
14:00
→
14:05
General news - Review of minutes 5mSpeakers: Erwin Siesling (CERN), Hanne Heylen (CERN), Joachim Vollaire (CERN)
-
14:05
→
14:15
Safety matters 10mSpeakers: Andreas Herty (CERN), Hanne Heylen (CERN), James Devine (CERN), Joachim Vollaire (CERN)
-
14:15
→
14:20
IBDRS project update - Work in the primary areas - BTY line @ 2 GeV 5mSpeakers: Ana-Paula Bernardes (CERN), Joachim Vollaire (CERN), Melania Averna (CERN)
-
14:20
→
14:25
RILIS Laboratory upgrade project - Exhaust gas line re-routing - activities update 5mSpeakers: Carlotta Porzio (CERN), Katerina Chrysalidis (CERN), Laura Lambert (CERN), Sebastian Rothe (CERN)
- Planning in the hall for Friday 6th? Is it possible to use the crane to bring in some pipes for HRS vac pipe rerouting?
- Shielding installation
- Shoring of HRS entry this friday 6th
- Installation of shielding in back wall of RILIS next week
- Shoring to be removed tentatively next fri 13th
- Back pillar being removed and modified today
- Roof I-beam also to be temporarily removed to allow for forklift to fit in with the crane
-
14:25
→
14:35
Machine related activities in the experimental hall 10mSpeakers: Erwin Siesling (CERN), Simon Mataguez (CERN)
FEs :
- YHRS.QP30-NEG issue - few tests carried out by L. Nies & E. Siesling from the HRS sep zone on the cables to the HRS FE: For the HRS.QP30-L & R there is a low resistance (few kOhms) at 1kV already and the HT tester indicates 'breakdown'. For the HRS.QP30-T & B things are fine with very high resistance and up to 5kV. The HRS10 sector was/is under vacuum.
- Cabeles to the HRS10 sector elements are all disconnected from the patch panel towards the FE in the HRS separator zone and should be reconnected if convertor tests (workingset to element) are to be carried out.
Platform EPC / HW tests:
- HW tests finishing: M. Lozano / S. Mataguez. Reporting will follow (M. Lozano).
- Polarity and capacitance tests carried out on all elements except for the HRS40 sector (TPs ISCOOL no water - hence no vacuum) and HRS10 and GPS10 with no access to do polarity tests (porposal E. Siesling - do the controls, ccv, aqn tests from the workingset. For this HRS10 sector elements need to be reconnected)
- Mon 16th Nov vacuum exhaust lines flushing, etc. - vacuum stop. Vacuum system fully consigned Mon 23rd. Rerouting of the (RILIS related) vacuum exhaust tubes - no VSC til May '27.
REX/HIE ISOLDE:On the List:
- REX/HIE LINAC:
- On the list: replacement prim. pumps CMs, backing & roughing lines - Oct/Nov (after LHC LS high peak work).
- Ongoing: Modification of the chamber of the XL9GP Quad. Chamber to be back in the machine next summer latest.
- Additional NEG pump is to be added onto the XLN2 DBox: After re-installation end of '27 - Replacement of the 9 GAP structure vacuum seal: Response from the RF group (S. Ramberger):
We would in a first step want to try to retighten the bolts.
Ongoing:
- Our analysis showed that we might not have applied the maximum torque that is acceptable.
- Only if the structure is still not leak tight, we would open it, verify the exact sealing dimensions and replace the seal with the one that we have available.
- If it would still not be leak tight, we would manufacture a new seal according to the measured sealing dimensions.
- we need to make sure that we can power the 9GAP after each action for testing. Remark E.S.: tests can only be done after the vacuum and water (when RF to the structure) are back (resp. May, July '27) and RF amplifiers are installed. - REX RF room:
- old systems are being dismounted (G. Piccinini) for removal which will be 3 racks on each chassis, the doors of the RF room needed to be removed (stored on the HIE tunnel roof)
- due to the size and weight the yellow passerelle will need reinforcement and enlargement by scaffolding: calculations ongoing, date of scaffolding installation tbc - users/dewar handling will coordinated. Blockage of the 'hole'from the yellow passerelle down is OK with REX if temporary (estimation 1 wk) but to kep in mind that this is the only way to exchange the REX solenoid if needed (F. Wenander BE/ABP). - REX Low E development:
- Dedicated meeting Fri 6 Nov (M. Nieto) to discuss in more detail the replacement of the leaking REX mass slits actuator for a bellow type one using a HIE type actuator. Under discussion/investigation M. Nieto/A. Rodriguez/E. Siesling/F. Wenanderfor the newer version. - REXTRAP innner structure work ongoing in 179. (J. Schoonwater, F. Wenander).
- BI work in the HEBT zone is to start after PUMA / RC6 related BI work has finished:
A dedicated meeting on the upcoming work organised (S. Mataguez) Thursd 5 Nov covering:LS3 Beam instrumentation improvements:- Installation of 3 new collimator slit blades (square stroke) replacing 3 standard slit blades (circular stroke)
- Installation of 3 new linear stepping motor actuators
- Replacement/reordering of 6 linear stepping motor actuators
- Recalibration of 13 beam diagnostic boxes
Separator zones HRS & GPS/HRS entrance area:
- Field regulation rack arrived (in SAS: M. Torre / E. Siesling). This will occupy the position of the ex-rack60 and could be pt in place before cable tray installation.
- Installation cable trays foreseen before the end of the year - J.P. Billon-Grand/RILIS. Dimensions Tapestation and Field regulation racks have been transfered to integration (V. Barozier)
- PLAN (14522) created for the field regulation installation works (M. Torre / E. Siesling) - follow up to start
- DCCT boxes (M. Marreiros EN/EL, D. Giloteaux TE/MSC, M. Torre TE/MSC) for HRS90 & 60 accurate DC current measurement for the field regulation systems positions to be validated by the take holders (S. Marzari, K, Chrysalidis, M. Duraffourg, V. Barozier).
- Thursday 5 Nov access HRS sep zone to pull back HRS field regulation cables (M. Torre)
- ISCOOL area: Boris tube reinforcement beams will be mounted in week 47.
Infrastructure / 197 BTY / Separators Power supplies consolidation:
- 197 installation CV tubing works ongoing: See comments J. Vollaire
- Chilled water / AC in 508 and for the 170 HVAC system situation: flexibles on the roof of 170 being installed. Not to delay the installation works the choise of material to be chosen will probably be the type holding halogene with a derogation from HSE (since it is on the roof). Cheaper solution and under thershold for tender (A. Fournier, J. Vollaire).
- For info: Discussion/study within EN/CV on a possible different routing of the cooling water piping so that it is diconnected from the IBDRS civil works interms of timeline for installation of a new pipe rack. This is to allocte sufficient resources for the design and installation in the view of having water back in the summer of '27 and more important to ensure the early restart of the facility in '28 (S. Deleval, N. Szczepanik-Scislo, G. Petrika, A. Fournier EN/CV)
Fume cupboards GLM/GHM:
- New chemical fumecupboard in old ICR: - POssibilities being investigated on how to rout the duct through 197 (cable tray power room) (N. Szczepanik-Scislo, A. Fournier, V. Barozier)
Fiber Optics work / ISOLDE timing upgrade:
- 2nd phase of the White Rabbit fiber optics installation (N. Simic) postponed to 14/06/27 - 25/06/27
- ISOLDE Central Timing Renovation - Functional and operational requirments being written by J. Sanchez
- Regular meetings (L. Nies, E. Piselli, G.P. Di Giovanni, Controls groups). This includes the CAO switching, REX timing upgarde, MR-Tof/PUMA, beamgate cons.
- ECR for beamgates will be rewritten (L. Nies).
- New beamgates power supplies will installed in the OICR rack 03: Agreed is to move the 3 FEC's - equipment owners been contacted (L. Nies). Ongoing.
Other:
- evacuation of the buildings in the frame of the installation of the large standing crane for IBDRS moved to 18th Nov (still valid?) (G. Cnudde). Confirmation will be send out in due to the ISO House group
-
14:35
→
14:40
Experiments related activities in the experimental hall 5mSpeakers: Hanne Heylen (CERN), Oscar Fjeld (CERN)
-
14:40
→
14:45
Experiments related activities in the experimental hall 5mSpeaker: Erwin Siesling (CERN)
ISOLTRAP:
- Re-installation of infrastructure and equipment ongoing (A. Jaries)
- AC cabling to be discussed with EN/EL P. Gallay and new contractor (L. Nies, A. Jaries, E. Siesling)
- Crane lock-out wk47: dewar handling planned for Fri 20 Nov - OK (see PUMA AC cablling works below)
COLLAPS/CRIS platform:
- Meetings held in situ with Tranport and SCE (G. Venel & team, M. Boubon, S. Grail (Ouvaroff), S. D'Armancourt) for the removal of the concrete structure:
First stage:
- wk 45/46: removal as much as possible of the equipment / mirror boxes, etc. - Users in situ 2 wks (2-13 Nov) (Liss, Kara, Jessica, Tim)
- wk 47 (probably 18/19 Nov): scaffolding for transport to prepare for the handling of the large beam (8m, 6 ton) + access to remove supports, laser guidance tubes, etc. (Ask O. Fjeld + users? + E. Siesling)
- wk 47/48 (as of 23 Nov): the large concrete beam will be removed + 2 small 80x80cm blocks (transport, G. Venel)
- wk 48 cutting up of the concrete beam in 3 pieces in the ISOLDE hall (PUMA area, wet cutting with plastic SAS) (SCE, S. D'Armancourt) - dates cutting to be confirmed
For this the area behind Miniball / PUMA needs clearing/moving of some items - action E. Siesling, H. Heylen, users
second stage:
- removal of the scaffodling WK48/49
- followed by removal of the services on the tall concrete blocks (RP monitors, cable trays, AC distribution). AC distribution removal (keep boxes) discussed with P. Gallay EN/EL - ticket probably enough
- removal of the tall concrete blocks
-timeline: before the end of the year - Organisational meeting held Tues 15h (H. Heylen et all.)
PUMA:
- EN/EL AC cable work VIC held last Friday:
- AC cable installation (diesel and network) wk47 (16-20 Nov) from 199 along the walls in the HIE side of the hall to a new LC7 AC cabinet.
- Crane will be locked out wk47: - ISOLTRAP dewar Fri 20th morning should work, Ok with REX team.
- extra I-pillar for the AC cabinet and small I-beam + hor beam for AC boxes will be placed by the contractor: Please keep an eye on position: L. Nies, M. Kirschbaum
- AC cable holders side of LC7 later (2 AC power cables: PUMA footprint length + 2 meters) - AUG control cable as part of the AC installation needs to be pulled from the PUMA AC cabinet (LR7): wk47/48 -> to be coordinated with COLLARIS scaffolding installation and 197 CV works (EN/EL, E. Siesling, A. Fournier, J. Vollaire)
ISS
- Transport would like to borrow the yellow ISS end shielding plates carrier for a few days (end this week till mid next week) for a handling actiovity at CNGS - ISS owners have been contacted (A. Dolan, L. Gaffney)
-
14:45
→
14:55
Mechanical Workshop and other support requests 10mSpeaker: Oscar Fjeld (CERN)
Completed work:
- VITO/DNP (193) - modified a stainless-steel tube and machined a groove in a laser-cut component.

- VITO/DNP (192) - drilled through two Teflon rods, retaining the greatest practical length.

- ASCII/GLM (189) - made a new cable opening in an acrylic sheet.
- GLM (188) - cut a PVC pipe and made an opening in a perforated sheet
- GPS/HRS (180) - drilled mounting holes for an exhaust bracket on each side of the wall.

- Test mount demount valve and pump frontend offline 2
Development and follow-up:
- Building 170 ISOLDE sign (88) - ordered.
New requests:
- Miniball (190) - design, manufacture and install supports for the new preamplifier boxes.
- Miniball (191) - refurbish the beamdump rack: dismantle the existing vacuum controls and fit new rails and Perspex panels.
-
14:55
→
15:00
Planned visits and AOBs 5mSpeaker: Hanne Heylen (CERN)
- Request from Guilherme:
- Painted reference lines on the floor - defined by theodolite team - for alignment of experiments, defining beam axis at the center of the beamline
- As long as possible (several meters) to cope with the diff sizes and mounting setups
- Proposal to have them at GLM, LA1, LA2, and some others TBD
- Request from Guilherme:
-
14:00
→
14:05

